# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ B.5 Proof of Proposition 3
$\mathbb{E}\left[\dot{m}_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}\right]=\mathbb{E}\left[z_{\theta}(X_{i},A_{i})Y_{i}+v_{\theta}(X_{i},A_{i})|X_{i},A_{i}\right]=z_{\theta}(X_{i},A_{i})\mathbb{E}[Y_{i}|X_{i},A_{i}]+v_{\theta}(X_{i},A_{i}).$
@@
$g_{t}(X_{i},a)-\mathbb{E}\left[m_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}\right]=z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)f_{t}(X_{i},a)+v_{\bar{\theta}_{T}}(X_{i},A_{i})-v_{\theta^{\star}}(X_{i},A_{i}).$
@@
$z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)E[Y_{i}|X_{i},A_{i}]$
$=z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)E[Y_{i}|X_{i},A_{i}]+z_{\bar{\theta}_{T}}(X_{i},a)E[Y_{i}|X_{i},A_{i}]-z_{\bar{\theta}_{T}}(X_{i},a)E[Y_{i}|X_{i},A_{i}]$
$=(z_{\bar{\theta}_{T}}(X_{i},a)-z_{\theta^{\star}}(X_{i},a))E[Y_{i}|X_{i},A_{i}]+z_{\bar{\theta}_{T}}(X_{i},a)\left(f_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}]\right).$
Applying the assumptions that $\bar{\theta}_{t}\overset{p}{\to}\theta^{\star}$ and $f_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}]\overset{p}{\to}0$ along with the continuous mapping theorem immediately yields the result that $g_{t}(X_{i},a)$ is consistent.
Next, we will prove the second part of the theorem. By the properties of variance and because $z_{\theta}(X_{i},A_{i})$ is fixed conditional on $X_{i},A_{i}$ we have that,
$\mathrm{Var}\left[\dot{m}_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}\right]=z_{\theta}(X_{i},A_{i})z_{\theta}(X_{i},A_{i})^{T}\mathrm{Var}(Y_{i}|X_{i},A_{i}).$
By exactly the same reasoning as above we have that
$h_{t}(X_{i},a)-\mathrm{Var}\left[\dot{m}_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}\right]$ $=(z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{T}-z_{\theta^{\star}}(X_{i},a))z_{\theta^{\star}}(X_{i},a)^{T})E[Y_{i}|X_{i},A_{i}]$
$+z_{\bar{\theta}_{T}}(X_{i},a)\left(e_{t}(X_{i},a)-E[Y_{i}^{2}|X_{i},A_{i}]\right)$
$+z_{\bar{\theta}_{T}}(X_{i},a)\left(f^{2}_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}]^{2}\right).$
Applying the assumptions that $\bar{\theta}_{t}\overset{p}{\to}\theta^{\star}$, $f_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}]\overset{p}{\to}0$, $e_{t}(X_{i},a)-E[Y_{i}^{2}|X_{i},A_{i}]\overset{p}{\to}0$ along with the continuous mapping theorem immediately yields the result that $h_{t}(X_{i},a)$ is consistent.
@@ B.6 Proof of Proposition 4
$\leq\mathbb{P}\Bigl(\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}R_{t}(\theta)-\mathbb{E}\left[R_{t}(\theta)|\mathcal{H}_{t-1}\right]\right\}+\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[R_{t}(\theta)-R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}$
$-\frac{1}{T}\sum_{t=1}^{T}\left\{R_{t}(\theta^{\star})-\mathbb{E}\left[R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}\geq 0\Bigr).$
@@
$\mathbb{E}\left[R_{t}(\theta)|\mathcal{H}_{t-1}\right]$ $=\mathbb{E}\left[\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(m_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{m_{\theta}(X_{t},A_{t},Y_{t})-m_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|\mathcal{H}_{t-1}\right]$
$=\mathbb{E}_{\mathcal{P},p_{t}}\left[\frac{\pi_{e}(A_{t}=a|X_{t})}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}m_{\theta}(X_{t},A_{t},Y_{t})\right]$
$\quad\quad+\sum_{a=1}^{k}\mathbb{E}_{\mathcal{P},p_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)m_{\theta}(a,X_{t},f_{t}(a,X_{t})\right]$
@@
Thereforore,
$\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[R_{t}(\theta)-R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}$ $=\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\mathbb{E}_{\pi_{e}}\left[m_{\theta}(X_{t},A_{t},Y_{t})-m_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\right\}<-\delta_{1}$
for some $\delta_{1}>0$ by 3. Finally, we apply Lemma 1 in conjunction with 5 to conclude that the first and third terms converge. Thus, $\mathbb{P}\left(\left\lVert\tilde{\theta}_{T}-\theta^{\star}\right\rVert\geq\epsilon\right)\rightarrow 0$.
@@ B.7 Proof of Proposition 5
By the WLLN, $\frac{rT}{n_{r}}\overset{p}{\to}1$. We also have that $\hat{\nu}_{t}-\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right]\overset{p}{\to}0$ by the same argument given in the proof to Proposition 2.
Let us rewrite the first term in the expression as
$\frac{rT}{n_{r}}\frac{1}{rT}\sum_{X_{i}\in\tilde{X}}\hat{\nu}_{t}(X_{t})\hat{\nu}_{t}(X_{t})^{T}-\left(\frac{rT}{n_{r}}\frac{1}{rT}\sum_{t=1}^{T}\hat{\nu}_{t}(X_{t})\right)\left(\frac{rT}{n_{r}}\frac{1}{rT}\sum_{t=1}^{T}\hat{\nu}_{t}(X_{t})\right)^{T}.$
By the continuous mapping theorem and Slutsky’s theorem, it is sufficient to show that
$\frac{1}{rT}\sum_{t=1}^{T}\zeta_{t}\hat{\nu}_{t}(X_{i})\overset{p}{\to}\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right];$
$\frac{1}{rT}\sum_{t=1}^{T}\zeta_{t}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}\overset{p}{\to}\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})\dot{m}_{\theta}(X_{t},A_{t},Y_{t})^{T}|X_{t}\right].$
Let us substitute $\hat{\nu}_{t}(X_{i})=\hat{\nu}_{t}(X_{i})-\nu^{\star}_{t}(X_{t})+\nu^{\star}_{t}(X_{t})$. We then have that
$\frac{1}{rT}\sum_{t=1}^{T}\zeta_{t}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}=\frac{1}{rT}\sum_{t=1}^{T}\zeta\nu^{\star}_{t}(X_{t})\nu^{\star}_{t}(X_{t})^{T}+o_{p}(1),$
$\frac{1}{rT}\sum_{t=1}^{T}\zeta_{t}\hat{\nu}_{t}(X_{i})=\frac{1}{rT}\sum_{t=1}^{T}\zeta\nu^{\star}_{t}(X_{t})+o_{p}(1).$
Both of these expressions converge in probability to the appropriate expectations by the WLLN.
For the second term, let us rewrite as
$\frac{1}{n_{r}}\sum_{t=1}^{T}\sum_{a\in\mathcal{A}}\zeta_{t}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}g_{t}(X_{i},a)=\frac{rT}{n_{r}}\frac{1}{rT}\sum_{t=1}^{T}\sum_{a\in\mathcal{A}}\zeta_{t}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}g_{t}(X_{i},a).$
By Slutsky’s Theorem, it is sufficient to show that
$\frac{1}{rT}\sum_{t=1}^{T}\sum_{a\in\mathcal{A}}\zeta_{t}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}g_{t}(X_{i},a)-\mathbb{E}\left[\sum_{a\in\mathcal{A}}\frac{\pi_{e}(A_{t}=a|X_{t})^{2}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\mathrm{Var}\left(\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))|X_{t}\right)|\mathcal{H}_{t-1}\right]=o_{p}(1).$
However, by independence of $\zeta_{t}$,
$\mathbb{E}\left[\frac{1}{rT}\sum_{t=1}^{T}\sum_{a\in\mathcal{A}}\zeta_{t}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}g_{t}(X_{i},a)\right]$ $=\frac{1}{r}\mathbb{E}\left[\zeta_{t}\right]\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\sum_{a\in\mathcal{A}}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1}^{(2)})}g_{t}(X_{i},a)\right]$
$=\mathbb{E}\left[\sum_{a\in\mathcal{A}}\frac{\pi_{e}(A_{t}=a|X_{t})^{2}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\mathrm{Var}\left(\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))|X_{t}\right)|\mathcal{H}_{t-1}\right].$
Therefore, this is true by the WLLN.
@@ Appendix C Procedure for Semi-Synthetic Data Construction and Additional Results
    $E[Y_{t}|X_{t},A_{t}]=\sum_{a=1}^{k}\beta^{a}_{1}\mathbbm{1}_{A_{t}=k}+\sum_{a=1}^{k}\beta_{2}^{a}f(X_{t})\times\mathbbm{1}_{A_{t}=k},$
@@
For the empirical simulations, we choose four different initializations
@@
    Scenario 1. Homoskedasticity with
    $\beta_{1}=(0,0,1,2,2,3,4,5),\beta_{2}=(1,2,3,4,5,5,5,5),\text{and unit variance uniformly}.$
    Results are shown in Figure 3.
@@
    Scenario 2 Heteroskedasticty across $X_{t}$ but not $A_{t}$
    $\beta_{1}=(0,0,1,2,2,3,4,5),\beta_{2}=(1,2,3,4,5,5,5,5),\text{and}\gamma=(1,1,1,1,1,1,1,1)\times 0.2.$
    Results are shown in Figure 4.
@@
    Scenario 3 Heteroskedasticty across $X_{t}$ and $A_{t}$ with
    $\beta_{1}=(0,0,1,2,2,3,4,5,5),\beta_{2}=(1,2,3,4,5,5,5,5),\text{and}\gamma=(1,2,3,4,5,5,5,5)\times 0.2.$
@@
    Scenario 4
    $\beta_{1}=(0,0,1,2,2,3,4,5,5),\beta_{2}=(0,0,0,0,0,0,0,0),\text{and unit variance uniformly}.$
@@
(a) Coverage for target $1-\alpha=0.8$
@@
Figure 3: Simulation results for scenario 1. The results are broadly aligned with those shown in Figure 2, except that using naive estimates of the covariance matrix for the MAIPWM estimate now does result in error control.
(a) Coverage for target $1-\alpha=0.8$
@@
Figure 4: Simulation results for scenario 1. This is a repeated view of the results shown in Figure 2. GLMs using naive inverse propensity weighting often undercover, especially in situations where the assignment probabilities vary substantially over time. We note that both sample splitting and external data reuse for covariance estimation are valid, but sample splitting has significantly wider confidence intervals. Reusing the same data for variance estimation and parameter estimation performs similarly to using external data, though we lack theoretical guarantees for this method.
(a) Coverage for target $1-\alpha=0.8$
@@
Figure 5: Simulation results for scenario 3, where additional heteroskedasticity is added across $X_{t}$ and $A_{t}$. The results are broadlyt similar to Figure 2, with the most obvious departure the MAIPWM (naive) method eventually reaches nominal coverage with sufficiently large amounts of samples.
(a) Coverage for target $1-\alpha=0.8$
@@
Figure 6: Simulation results for scenario 4, where the model is correctly specified. In these settings, we see that all methods cover, but it often takes the naive approaches (IPW and MAIPWM – Naive) estimates significantly more samples to reach the nominal coverage rates.

### v2

@@ B.5 Proof of Proposition 3
$\mathbb{E}\left[\dot{m}_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}=a\right]=\mathbb{E}\left[z_{\theta}(X_{i},a)Y_{i}+v_{\theta}(X_{i},a)|X_{i},A_{i}=a\right]=z_{\theta}(X_{i},a)\mathbb{E}[Y_{i}|X_{i},A_{i}=a]+v_{\theta}(X_{i},a).$
@@
$g_{t}(X_{i},a)-\mathbb{E}\left[m_{\theta^{\star}}(X_{i},A_{i},Y_{i})|X_{i},A_{i}=a\right]=z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)E[Y_{i}|X_{i},A_{i}=a]+v_{\bar{\theta}_{T}}(X_{i},a)-v_{\theta^{\star}}(X_{i},a).$
@@
$z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)E[Y_{i}|X_{i},A_{i}=a]$
$=z_{\bar{\theta}_{T}}(X_{i},a)f_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)E[Y_{i}|X_{i},A_{i}=a]+z_{\bar{\theta}_{T}}(X_{i},a)E[Y_{i}|X_{i},A_{i}=a]-z_{\bar{\theta}_{T}}(X_{i},a)E[Y_{i}|X_{i},A_{i}=a]$
$=(z_{\bar{\theta}_{T}}(X_{i},a)-z_{\theta^{\star}}(X_{i},a))E[Y_{i}|X_{i},A_{i}=a]+z_{\bar{\theta}_{T}}(X_{i},a)\left(f_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}=a]\right).$
Applying the assumptions that $\bar{\theta}_{t}\overset{p}{\to}\theta^{\star}$ and $f_{t}(X_{i},a)-E[Y_{i}|X_{i},A_{i}=a]\overset{p}{\to}0$ along with the continuous mapping theorem immediately yields the result that $g_{t}(X_{i},a)$ is consistent.
Next, we will prove the second part of the theorem. By the properties of variance and because $z_{\theta}(X_{i},A_{i})$ is fixed conditional on $X_{i}$ and $A_{i}$, we have that
$\mathrm{Var}\left[\dot{m}_{\theta}(X_{i},A_{i},Y_{i})|X_{i},A_{i}=a\right]=z_{\theta}(X_{i},a)z_{\theta}(X_{i},a)^{T}\mathrm{Var}(Y_{i}|X_{i},A_{i}=a).$
Analyzing the term,
$h_{t}(X_{i},a)$ $-\mathrm{Var}\big(\dot{m}_{\theta^{\star}}(X_{i},A_{i},Y_{i})\mid X_{i},A_{i}\big)$
$=z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}j_{t}(X_{i},a)-z_{\theta^{\star}}(X_{i},a)z_{\theta^{\star}}(X_{i},a)^{\top}\mathrm{Var}(Y_{i}|X_{i},A_{i}=a)$
$=z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}j_{t}(X_{i},a)-z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}\mathrm{Var}(Y_{i}|X_{i},A_{i}=a)$
$\quad\quad+z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}\mathrm{Var}(Y_{i}|X_{i},A_{i}=a)-z_{\theta^{\star}}(X_{i},a)z_{\theta^{\star}}(X_{i},a)^{\top}\mathrm{Var}(Y_{i}|X_{i},A_{i}=a)$
$=z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}\left(j_{t}(X_{i},a)-\mathrm{Var}(Y_{i}|X_{i},A_{i}=a)\right)$
$\quad\quad+\left(z_{\bar{\theta}_{T}}(X_{i},a)z_{\bar{\theta}_{T}}(X_{i},a)^{\top}-z_{\theta^{\star}}(X_{i},a)z_{\theta^{\star}}(X_{i},a)^{\top}\right)\mathrm{Var}(Y_{i}|X_{i},A_{i}=a).$
Applying the assumptions that $\bar{\theta}_{t}\overset{p}{\to}\theta^{\star}$, $j_{t}(X_{i},a)-\text{Var}[Y_{i}|X_{i},A_{i}=a]\overset{p}{\to}0$ together with the continuous mapping theorem immediately yields the result that $h_{t}(X_{i},a)$ is consistent.
@@ B.6 Proof of Proposition 4
$\leq\mathbb{P}\Bigg(\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}R_{t}(\theta)-\mathbb{E}\left[R_{t}(\theta)|\mathcal{H}_{t-1}\right]\right\}+\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[R_{t}(\theta)-R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}$-
$\qquad\qquad\frac{1}{T}\sum_{t=1}^{T}\left\{R_{t}(\theta^{\star})-\mathbb{E}\left[R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}\geq 0\Bigg).$
@@
$\mathbb{E}\left[R_{t}(\theta)|\mathcal{H}_{t-1}\right]$ $=\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(m_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{m_{\theta}(X_{t},A_{t},Y_{t})-m_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)|\mathcal{H}_{t-1}\right]$
$=\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\frac{\pi_{e}(A_{t}=a|X_{t})}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}m_{\theta}(X_{t},A_{t},Y_{t})\right]+$
$\qquad\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\mathbb{E}_{\mathcal{P},\pi_{t}}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)m_{\theta}(a,X_{t},f_{t}(a,X_{t}))\right]$
@@
Therefore,
$\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[R_{t}(\theta)-R_{t}(\theta^{\star})|\mathcal{H}_{t-1}\right]\right\}$ $=\sup_{\left\lVert\theta-\theta^{\star}\right\rVert\geq\epsilon}\left\{\mathbb{E}_{\pi_{e}}\left[m_{\theta}(X_{t},A_{t},Y_{t})-m_{\theta^{\star}}(X_{t},A_{t},Y_{t})\right]\right\}<-\delta_{1},$
for some $\delta_{1}>0$ by 3. Note that 5 implies that the bracketing number $N_{[]}(\epsilon,\mathcal{M}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$ for $\mathcal{M}_{\Theta}=\{m_{\theta}(X_{t},A_{t},Y_{t}):\theta\in\Theta\}$. This combined with 6 allows us to apply Lemma 1 conclude that the first and third terms also conerge to 0. Therefore, $\mathbb{P}\left(\left\lVert\tilde{\theta}_{T}-\theta^{\star}\right\rVert_{1}\geq\epsilon\right)\rightarrow 0$.
@@ B.7 Proof of Proposition 5
By the law of large numbers, we have that $\frac{n_{p}}{T}\rightarrow{p}p$ and therefore $\frac{1}{n_{p}}=\frac{1}{pT}+o_{p}(T^{-1})$. Define the per-round variance summand
$\Psi_{t}:=\Big(\widehat{\nu}_{t}(X_{t})-\overline{\nu}_{t}\Big)\Big(\widehat{\nu}_{t}(X_{t})-\overline{\nu}_{t}\Big)^{\top}+\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}\,h_{t}(X_{t},a)\;-\;\widehat{\nu}_{t}(X_{t})\widehat{\nu}_{t}(X_{t})^{\top}.$
We can therefore write $\widehat{V}_{t}=\frac{1}{n_{p}}\sum_{t=1}^{T}\zeta_{t}\,\Psi_{t}$. Moreover, $\widehat{V}_{t}=\frac{1}{pT}\sum_{t=1}^{T}\zeta_{t}\,\Psi_{t}+o_{p}(1)$.
By assumption, $\{\zeta_{t}\}$ are independent of all histories, therefore $\mathbb{E}[\zeta_{t}\Psi_{t}\mid\mathcal{H}_{t-1}]=p\,\mathbb{E}[\Psi_{t}\mid\mathcal{H}_{t-1}]$. Consequently, by the law of large numbers and Slutsky’s theorem, $\widehat{V}_{t}=\mathbb{E}[\Psi_{t}\mid\mathcal{H}_{t-1}]+o_{p}(1).$ The remainder of the argument is identical to Proposition 2 and, therefore,
$\big\|\widehat{V}_{t}-V_{t,\theta^{\star}}\big\|_{\mathrm{op}}\xrightarrow{p}0,$
as claimed.
@@ Appendix C Procedure for Semi-Synthetic Data Construction and Additional Results
    $E[Y_{t}|X_{t},A_{t}]=\sum_{a=1}^{K}\beta^{a}_{1}\mathbbm{1}_{A_{t}=a}+\sum_{a=1}^{K}\beta_{2}^{a}f(X_{t})\times\mathbbm{1}_{A_{t}=a},$
@@
For the empirical simulations, we choose six different initializations
@@
    Scenario 1: Correct Specification, Homoscedasticity, Unique Optimal Arm
    $\beta_{1}=(0,1,2,3,4,5,6,7),\beta_{2}=(0,0,0,0,0,0,0,0),\text{and unit variance uniformly}.$
    In this scenario, all methodologies other than naive unweighted MLE should perform relatively well as bandit algorithms will concentrate. Results are shown in Figure 3.
@@
    Scenario 2: Correct Specification, Homoscedasticity, No Unique Optimal Arm.
    $\beta_{1}=(0,0,1,2,2,3,5,5),\beta_{2}=(0,0,0,0,0,0,0,0),\text{and unit variance uniformly}.$
    In this scenario, IPW-style estimators may have difficulties but confidence intervals constructed as in Zhang et al. (2021) should still cover correctly. Results are shown in Figure 4.
@@
    Scenario 3: Misspecification, Homoscedasticity.
    $\beta_{1}=(0,0,1,2,2,3,4,4),\beta_{2}=(1,-1,1,0,1,1,1,-3),\text{and unit variance uniformly}.$
@@
    Scenario 4: Misspecification, Heteroskedasticity
    $\beta_{1}=(0,0,1,2,2,3,4,5),\beta_{2}=(1,-1,1,0,1,1,1,-2),\text{and }\gamma=(1,2,3,4,5,5,5,5)\times 0.2.$
@@
(a) Coverage for target $1-\alpha=0.9$
@@
Figure 3: Simulation results for scenario 1, where the model is correctly specified. In these settings, we see that all methods cover, but it often takes the naive approaches (IPW and SQ-IPW) estimates significantly more samples to reach the nominal coverage rates. Thompson sampling undercovers for several methods when propensity scores are not clipped, consistent with theory.
(a) Coverage for target $1-\alpha=0.9$
@@
Figure 4: Simulation results for scenario 2, where the model is correctly specified but there is not a unique optimal arm. The results are broadly similar as in Figure 4, though we note there is no theoretical guarantee that IPW estimates should cover in this setting (though they do empirically).
(a) Coverage for target $1-\alpha=0.9$
@@
Figure 5: Simulation results for scenario 3, when there is model misspecification but homoskedastic errors. IPW and SQ-IPW estimators now do not cover in every scenario, as their theoretical guarantees depend on correct specification of the working models.
(a) Coverage for target $1-\alpha=0.9$
@@
Figure 6: Simulation results for scenario 3, when there is model misspecification and heteroskedastic errors. The results are broadly similar to Figure 5.
Acknowledgments
Data and/or research tools used in the preparation of this manuscript were obtained and analyzed from the controlled access data sets distributed by the Osteoarthritis Initiative (OAI), a data repository located within the NIMH Data Archive (NDA). OAI is a collaborative informatics system created by the National Institute of Mental Health and the National Institute of Arthritis, Musculoskeletal and Skin Diseases (NIAMS) to provide a worldwide resource to quicken the pace of biomarker identification, scientific investigation and OA drug development. Dataset identifier(s): [outcomes99, kxrsqbu06, kxrsqbu00, enrollees, allclinical00, allclinical06].